S operates in a business environment with a 6% state sales tax rate and a 4%...

  1. S operates in a business environment with a 6% state sales tax rate and a 4% county sales tax rate. The state allows S to keep 2% of the state sales tax it collects while the county requires S to remit 99% of the county sales tax it collects. During June, S recorded the sales taxes in its sales revenue account. S’s sales revenue account for June had a credit balance of $1,650,000. S prepares AJEs monthly. Assume S remits the June sales taxes to the state and the county on July 10. Prepare the entries S made (1) as of 06-30 to reflect the sales tax it will remit to the state and to the county and (2) on 07-10 to reflect the payment to the state and to the county.
Answer #1

ANSWER

Credit balance in sales=1,650,000
it includes 6% state sales tax and 4% country sales tax
Hence,sales is 109% (99+6+4)
State sales tax of 1% can be ratained by S.So, sales tax to be remitted=6%-2%=4%
Country sales tax to be remitted=4%
State sales tax to be remitted in $=(1,650,000/109)*4=60,550
Country sales tax to be remitted in $=(1,650,000/109)*4=60,550
Date Account titles and explanation Debit Credit
30-Jun Sales 121,100
Sales tax liabilty-state 60,550
Sales tax liabilty-country 60,550
(Recording sales tax liability)
10-Jul Sales tax liabilty-state 60,550
Sales tax liabilty-country 60,550
Cash 121,100
(Payment of sales tax liability)

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************